v Analysis of Receivables Method At the end of the current year, Accounts Receivable has a...

v

Analysis of Receivables Method

At the end of the current year, Accounts Receivable has a balance of $410,000; Allowance for Doubtful Accounts has a debit balance of $3,500; and sales for the year total $1,850,000. Using the aging method, the balance of Allowance for Doubtful Accounts is estimated as $17,200.

a. Determine the amount of the adjusting entry for uncollectible accounts.
$

b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Bad Debt Expense.

Accounts Receivable $
Allowance for Doubtful Accounts $
Bad Debt Expense $

c. Determine the net realizable value of accounts receivable.
$
Answer #1

a) Adjusting entry amount = 17200+3500 = $20700

b) Determine balances

Account receivable 410000
Allowance for Doubtful Accounts 17200
Bad debt expense 20700

c) Net realizable value = 410000-17200 = $392800